*Again-Rules for rounding: If you are calculating in tablets, always round to the nearest 1/2 tablet....

  1. *Again-Rules for rounding:
    • If you are calculating in tablets, always round to the nearest 1/2 tablet.
    • When you are calculating IV' rates, round to the nearest whole number.
    • Perform rounding only the end of the calculation
  2. CONVERSTIONS TO USE
    1 grain =60 mg
    15 grain=1 gram
    1 pint= 2 cups=480mL
    1 quart= 2 pt= 4 cups=960mL
    1 oz= 30 mL
    1 cup = 8 oz = 240 mL
    1 quart = 2 pt = 4cups= 960 mL
    1cc= 1mL
    1tsp= 5mL
    1 Tbs= 15 mL
    1 gallon= 4 qt = 8 pt
  3. Note:Some drug conversions will use cc(cubic centimeters) This is the same as a mL. 1cc=1mL

(*There are actually 29.57 mL in 1 fl oz, but 30 mL is usually used by nurses in medical facilities. We will stay with this standard for all liquid measurements. 1 oz=30mL, 1 c=240mL, 1 pt=480mL, 1 quart=960mL.)

QUESTIONS:

1. A client receiving a prescription for LR 1,000 mL in 12hr. Using a micro drip system that delivers 60 micro drops/mL, the nurse will regulate the infusion for how many drops per minute.

_____gtts/min. (Round to the nearest whole number-remember no trailing zeros if a whole number)

2. You have an order to administer 1L of 5% Dextrose with 40 meq of potassium over 16 hours. How many mL per hour will you set on the IV pump?___________mL/hr.

(Remember to Round to the nearest whole number)

3. The provider has ordered 1,000 mL of 0.9% sodium chloride (normal saline) to infuse over 8 hr. You have tubing with a drop factor of 15 gtt/mL. You must calculate how many gtt/min to use to set the IV flow rate._______gtts/min (round to the nearest whole number)

4.Critical considerations for prevention of medication errors include all except

a. Follow the 4 rights of safe drug administration

b. None of these are true

c.All of these are correct.

d. Always check drug calculations with a coworker

e. Watch for look alike sound alike

f. Be knowledgeable about the purpose, action side effects, abnormal reactions, delivery methods and follow up
Answer #1

1)

Drop per minute= Total volume×Drop factor/Time in minute

Total volume= 1000ml

Time= 12Hr = 720 minutes

Drop factor= 60

So 1000×60/720= 83.33 Near round is

83 gtt. per minutes

2)

Ml per hour= Total volume/Number of hours

Total volume= 1000 ml

Time= 16 hr

1000/16= 62.5ml/hr

3)

Drop per minute= total volume× Drop factor/ time in minutes

Total volume= 1000 ml

Drop factor= 15 gtt. Per ml

Time= 8hr= 480 minutes

1000×15/480= 31.25 round near is

31gtt. Per minutes